# DeltaPhi: Learning Physical Trajectory Residual for PDE Solving

### v1

@@ Abstract
Although neural operator networks theoretically approximate any operator mapping, the limited generalization capability prevents them from learning correct physical dynamics when potential data biases exist, particularly in the practical PDE solving scenario where the available data amount is restricted or the resolution is extremely low. To address this issue, we propose and formulate the Physical Trajectory Residual Learning (DeltaPhi), which learns to predict the physical residuals between the pending solved trajectory and a known similar auxiliary trajectory. First, we transform the direct operator mapping between input-output function fields in original training data to residual operator mapping between input function pairs and output function residuals. Next, we learn the surrogate model for the residual operator mapping based on existing neural operator networks. Additionally, we design helpful customized auxiliary inputs for efficient optimization. Through extensive experiments, we conclude that, compared to direct learning, physical residual learning is preferred for PDE solving.
@@ 1 Introduction
Due to lacking analytical solutions, solving complex Partial Differential Equations (PDEs), e.g. Navier-Stokes, relies on numerical simulation based on ultra-fine grid division, which consumes expensive computational resources and time. To resolve this, machine learning based PDE solving [1, 2, 3, 4, 5] has sparked extensive research interest in recent years. Notably, the operator learning [6, 2, 1] directly learns the mapping between the input-output function fields, forming a surrogate model for a given parametric PDE. Then the solution for any input functions could be quickly obtained via neural network forward computation.
Despite current operator learning networks [2, 3] theoretically being able to approximate any operator mapping, their generalization capabilities remain limited. First, the networks are prone to overfitting to training resolution while suffering performance degradation on other resolutions [7], especially when the training resolution is extremely low. Second, when available training data is limited, the network may struggle to accurately model the correct physical dynamics due to potential biases in the data. Third, data scarcity and bias are common occurrences in practical PDE solving [8, 9]. Therefore, further investigation is required to enhance the neural operator generalization ability.
In this work, for generalized PDE solving, we introduce the Physical Trajectory Residual Learning (DeltaPhi), which reformulates the direct operator mapping between the input-output function fields as the mapping between input function pairs and output function residuals (residual operator mapping). The residual operator mapping could be learned with any existing neural operators (such as FNO [2], NORM [10], FFNO [11], etc. ), thereby obtaining corresponding residual neural operator. First, we reconstruct the training set from data pairs of direct operator mapping to data pairs of residual operator mapping, via assigning an auxiliary trajectory to every training data. The auxiliary trajectory is a different trajectory sampled from the training set based on similar trajectory retrieval. Next, we optimize the neural operator network with reconstructed training data. By concatenating the pending solved input function and auxiliary trajectory together as input of the network, and adding the auxiliary trajectory solution to the final output of the network via residual connection, we could optimize the network with existing training frameworks. All other setups (including the loss function, trainable weights amount, etc. ) remain consistent with the base operator networks. Compared to previous direct operator learning, the reconstructed residual training labels could be customized with enhanced diversity through selecting appropriate auxiliary samples. Consequently, the restricted diversity and potential distribution bias due to limited training data amount could be mitigated, thereby obtaining more generalized surrogate models.
Residual learning methodology has been extensively studied in the machine learning community [12, 13, 14, 15]. Previous works present various benefits derived from residual learning, such as enhancing generalization capability, handling complex regression mapping, etc. In this work, we introduce this concept into the neural operator learning task, a more complex regression task involving learning the mapping between high-dimension (tens of thousands) fields. The residuals between different but similar physical trajectories are learned, and utilized to solve PDEs. Through the visualization analysis of label distribution, we statistically validate the favorable generalization property of learning the residual operator mapping.
The contributions are summarized as follows:
  • We propose and formulate a novel framework for data-driven PDE solving in restricted data (limited amount, low resolution, biased distribution) scenarios, i.e. the Physical Trajectory Residual Learning (DeltaPhi).
  • We statistically analysis the advantage of residual operator learning and make instantiations based on similar trajectory retrieval and existing neural operator networks.
  • We validate the superiority of residual operator learning through extensive experiments across different PDE problems, different physical domains (both regular and irregular problems), different base models, different training set sizes, and different training resolutions. Additionally, detailed ablation studies are conduct.
2 Related Work
@@
Neural operator learning. Fourier Neural Operator (FNO) [2] utilizes the Fourier Transform based integral operation to efficiently learn the physical dynamics using limited training data. Factorized Fourier Neural Operator (FFNO) [11] factorize the Fourier Transform along each dimension, reducing the number of network weights. Clifford Fourier Neural Operator (CFNO) [16] employs the Clifford Algebra in the network architecture, incorporating geometry prior between multi physical fields. The various neural operators could be simply integrated with the proposed Physical Trajectory Residual Learning.
@@
Other works investigate architecture improvements for different needs, including chaotic systems modeling [17], physical-informed instance-wise finetuning [7], accelerating computation [18, 19, 20], spherical fields processing [21], irregular fields learning [22, 23], non-periodic boundary fields modeling [24], large-scale pretraining [25, 26], etc. In addition, some works explore other network backbones, e.g. improved frequency-spatial domain transformation [27, 28, 29, 30], convolutions [31], graph neural network [32, 6, 33], attention mechanism [34, 23, 35], and diffusion models [36]. Incorporating residual learning with these designs is worth exploring in the future.
@@
Residual learning in previous works. Some works learn residual between different data samples. Similar to ResMem [12], we also utilize the training set as auxiliary memory during inference, enjoying its generalization enhancement property. The difference lies that (1) we memorize the auxiliary labels and learn the residuals, while ResMem memorizes the residuals and learns the labels, (2) our framework is end-to-end trained while ResMem is trained in separate stages. Similar external memory augmented strategies are popular in the Language Modeling community [37, 38].
@@
Other works learn residuals between low resolution and high resolution representation for image super-resolution [39] and deep learning based CFD simulation [39]. Similar residual learning is investigated in time series prediction task  [40, 13]. The primary distinction between these works to ours is that we do not utilize the corresponding low-resolution solutions or previous time-step solutions, but instead, learn the residuals between different trajectories.
@@ 3 Methodology
3.1 Preliminary
3.1.1 Direct learning formulation
Direct operator mapping. Previous works [2, 11, 23] formulate the parametric PDE numerical solution as an operator mapping between two infinite dimensional function spaces:
$\mathcal{G}:\mathcal{A}\longrightarrow\mathcal{U},$ (1)
where $\mathcal{A}$ and $\mathcal{U}$ are both banach spaces of functions. For steady-state solving problems, e.g. Darcy Flow, $a(x)\in\mathcal{A}$ represents the diffusion coefficient function, and $u(x)\in\mathcal{U}$ is the corresponding solution function. For time-series forecasting problems, e.g. Navier-Stokes, $a(x)\in\mathcal{A}$ typically denotes the state function of the preceding few steps, while $u(x)\in\mathcal{U}$ represents the function in the subsequent few steps. The learning task is to study a parameterized surrogate model $\mathcal{G}_{\theta}$ for $\mathcal{G}$.
Domain discretization. Following [2, 41], we numerically process the function $a(x)$ and $u(x)$ by taking the discretization of the continuous domain. For regular domain problems, we uniformly discretize the regular field, using standard rectangular grids to capture the physical states. For irregular domain problems, we take the discrete mesh representation of the field.
Optimization problem. Consider the training set $\mathcal{T}$ consisting of $N$ i.i.d. pairs $\{a_{i},u_{i}\}_{i=1}^{N}$, where $a_{i}\in\mathcal{A}$, $u_{i}\in\mathcal{U}$, the objective is to find $\theta$ that minimizes the following cost function:
$\ell=\sum_{i=1}^{N}\mathcal{L}(\mathcal{G}_{\theta}(a_{i}),u_{i}),$ (2)
where $\mathcal{L}$ is a cost function that measures the difference between two functions. During inference, given the input function $a\in\mathcal{A}$, we obtain the corresponding solution $u=\mathcal{G}_{\theta}(a)$.
3.1.2 Neural Operator Network
Prior works [2, 34, 11] have extensively explored neural network architectures for learning the operator mapping $\mathcal{G}_{\theta}$ between function spaces. Among these, the Fourier Neural Operator (FNO)  [2, 1] has been the most commonly applied model across various physical systems. Neural operators are defined as multi-layer networks formulated as follows:
$\mathcal{G}_{\theta}=Q\circ\sigma(W_{l}+\mathcal{K}_{l})\circ\cdots\circ\sigma(W_{i}+\mathcal{K}_{i})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1})\circ P,$ (3)
where $l$ represents the number of layers in the network. The projector $P$ encodes the input function $a_{i}$ into an implicit embedding $v_{0}$. The operators $W_{i}$ and $\mathcal{K}_{i}$ at the $i$-th layer transform the input function $v_{i-1}$ into $v_{i}$. $\sigma$ is an element-wise non-linear operation introduced for learning non-linear mapping between function spaces. $Q$ is the output projector which decodes the last layer’s output $v_{l}(x)$ into the targeted function $u_{i}(x)$.
The components of $\mathcal{G}_{\theta}$ could be differently instantiated across various models. Typically, $P$ and $Q$ are implemented as fully connected layers. In Fourier Neural Operator [2], $\mathcal{K}_{i}$ is instantiated as an integral operator based on Fourier Transform, and $W_{i}$ is a linear fully connected layer. In NORM [10], $\mathcal{K}_{i}$ is implemented based on the Laplacian Transform for processing irregular domains. In addition, $\mathcal{K}_{i}$ can also be a convolution operation, adopting an architecture similar to Resnet [42].
Although direct operator learning networks can theoretically approximate any operator mapping, they encounter limited generalization capabilities in practical applications. First, the networks tend to overfit on trained resolution, and suffer performance drops on high resolution inference. Second, the model is vulnerable to potential bias in training data (especially when the data amount is limited), thus failing to capture the intrinsic physical trajectory dynamics.
3.2 Physical Trajectory Residual Learning
@@
Figure 1: The overall architecture of Physical Trajectory Residual Learning. Given an input trajectory $a_{i}$, we first sample a similar auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$ from the training trajectory set $\mathcal{T}$. Subsequently, the trajectory $a_{i}$ and $(a_{k_{i}},u_{k_{i}})$ are concatenated and fed into the the neural operator $\mathcal{G}_{\theta}^{\Delta}$, producing the predicted trajectory residual. Finally, the predicted solution $\hat{u}_{i}$ is obtained by adding the predicted trajectory residual with the auxiliary solution $u_{k_{i}}$.
3.2.1 Formulation
@@
Different from the aforementioned direct operator learning, we propose to learn the residual operator mapping that captures the differences between pairs of physical trajectories:
@@
$\mathcal{A}^{2}$ :$:\{(a_{i},a_{j})\mid a_{i}\in\mathcal{A},a_{j}\in\mathcal{A}\},$
$\Delta\,\mathcal{U}$ :$:\{u_{i}-u_{j}\mid u_{i}\in\mathcal{U},u_{j}\in\mathcal{U}\},$
@@
The residual mapping operator $\mathcal{G}^{\Delta}$ maps the input functions pair $(a_{i},a_{j})$ into the corresponding function residual $u_{i}-u_{j}$, i.e. $u_{i}-u_{j}=\mathcal{G}^{\Delta}(a_{i},a_{j})$. This approach leverages the concept of residuals to simplify the learning task. Instead of learning the direct operator mapping from $\mathcal{A}$ to $\mathcal{U}$, we focus on learning the differences between output functions. The transformed training label distribution, i.e. the function residual distribution, could have sufficient diversity by selecting appropriate auxiliary samples, thereby alleviating the training data bias and better handling the unknown function to be solved. We detail the explanation about this in Section 3.2.5.
Both direct operator mapping and residual operator mapping are mappings between Banach spaces of functions. Therefore, we could employ the off-the-shelf neural operator networks [2, 11] as the base network backbone. Neural operator networks are designed to handle mapping between function spaces and are known for their universal approximation capabilities, making them well-suited for residual operator mapping.
3.2.2 Auxiliary Trajectory Sampling
Given an input function (or named as input trajectory) $a_{i}$, we need an auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$ to learn the residual neural operator $\mathcal{G}^{\Delta}$ and infer the corresponding output $u_{i}$ of $a_{i}$. The auxiliary trajectory can be sampled from the existing data, allowing us to leverage prior knowledge to improve the generalization capabilities of the model. By selecting an appropriate auxiliary trajectory, we can simplify the computation and enhance the robustness of the residual operator mapping.
Given the training set $\mathcal{T}=\{a_{i},u_{i}\}_{i=1}^{N}$, each data pair $(a_{i},u_{i})$ is associated with an auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$, sampled from $\mathcal{T}$. The index $k_{i}$ of the auxiliary trajectory is defined by:
$k_{i}=\mathcal{F}^{ret}(a_{i},\mathcal{T}),\,k_{i}\in\{1,2,...,N\},\,k_{i}\neq i,$ (5)
where $\mathcal{F}^{ret}$ is the sampling function. A straightforward implementation of $\mathcal{F}^{ret}$ could be random sampling from $\{1,2,...,i-1,i+1,...N\}$. Using the calculated $k_{i}$, we transform the origin training set $\mathcal{T}$ to a residual training set $\mathcal{T}^{res}=\{(a_{i},a_{k_{i}}),u_{i}-u_{k_{i}}\}_{i=1}^{N}$. Below we present our implementation of $\mathcal{F}^{ret}$ in detail.
Training stage. During training, give an input function $a_{i}$, we randomly sample a trajectory $a_{j}$ from the set of trajectories with top-k similarity scores to $a_{i}$. The sampling function $\mathcal{F}^{ret}$ is defined as:
$\mathcal{F}^{ret}(a_{i},\mathcal{T})$ $=\texttt{RandSample}(\texttt{argtopK}_{j,j\neq i}\,\texttt{Sim}(a_{i},a_{j})),$ (6)
$\texttt{Sim}(a_{i},a_{j})$ $=1-{a_{i}\cdot a_{j}\over\|a_{i}\|\|a_{j}\|},$

### v2

# DeltaPhi: Physical States Residual Learning for Neural Operators in Data-Limited PDE Solving

@@ Abstract
The limited availability of high-quality training data poses a major obstacle in data-driven PDE solving, where expensive data collection and resolution constraints severely impact the ability of neural operator networks to learn and generalize the underlying physical system. To address this challenge, we propose DeltaPhi, a novel learning framework that transforms the PDE solving task from learning direct input-output mappings to learning the residuals between similar physical states, a fundamentally different approach to neural operator learning. This reformulation provides implicit data augmentation by exploiting the inherent stability of physical systems where closer initial states lead to closer evolution trajectories. DeltaPhi is architecture-agnostic and can be seamlessly integrated with existing neural operators to enhance their performance. Extensive experiments demonstrate consistent and significant improvements across diverse physical systems including regular and irregular domains, different neural architectures, multiple training data amount, and cross-resolution scenarios, confirming its effectiveness as a general enhancement for neural operators in data-limited PDE solving.
@@ 1 Introduction
Due to the lack of analytical solutions, solving complex Partial Differential Equations (PDEs) e.g. Navier-Stokes traditionally relies on numerical simulations based on ultra-fine grid division, which consumes expensive computational resources and time. This computational bottleneck has motivated the development of machine learning based PDE solving approaches (kovachki2021neural, ; li2020fourier, ; lu2019deeponet, ; raissi2019physics, ; long2018pde, ). In particular, neural operator learning (kovachki2021neural, ; li2020fourier, ; li2020neural, ) has emerged as a promising direction by directly learning the mapping between input-output function fields, enabling efficient solution prediction through neural network forward computation.
A critical challenge in neural operators is the severe scarcity of high-quality training data (bonnet2022airfrans, ; benitez2024out, ), which manifests in multiple aspects: (1) collecting comprehensive training data often require expensive physical experiments or high-resolution numerical simulations; (2) computational and storage constraints may restrict the resolution of available data; (3) the collected data frequently exhibits distribution bias due to physical or experimental limitations. These data limitations fundamentally constrain the generalization capability of current neural operators. To address these challenges, we propose DeltaPhi, a novel framework that helps neural operators learn from scarce data.
The core innovation of DeltaPhi lies in reformulating the PDE solving task from learning direct input-output mappings to learning the residuals between similar physical states. This reformulation enables implicit data augmentation by leveraging a key property of physical systems: states with similar initial conditions tend to follow similar solutions. By learning residuals between pairs of similar trajectories, DeltaPhi effectively expands the diversity of training samples without requiring additional data collection. This is fundamentally different from existing approaches that attempt to learn the straight mapping for each sample independently.
The proposed framework presents two advantages: (1) Enhanced Training Sample Diversity: The framework enables enhanced diversity in training labels through random auxiliary solution sampling, effectively mitigating distribution bias issues. This training distribution augmentation preserves physical validity while expanding the learned solution space. (2) Architecture Flexibility: DeltaPhi is architecture-agnostic and can seamlessly enhance existing neural operators (li2020fourier, ; chen2023learning, ; tran2021factorized, ) through a unified residual learning mechanism, making it applicable across different PDE solving scenarios.
We conduct extensive experiments to validate the effectiveness of physical states residual learning across various settings: (1) diverse physical systems including both regular domains (Darcy Flow, Navier-Stokes) and irregular domains (Heat Transfer, Blood Flow); (2) different neural architectures including spectral-based and attention-based operators; (3) varied numbers of training samples to test robustness under data scarcity; and (4) cross-resolution scenarios to evaluate generalization capability. The results consistently demonstrate that DeltaPhi significantly improves the performance of base models, particularly in data-limited settings.
Overall, this work introduces a promising framework for enhancing neural operators in practical PDE solving applications where high-quality training data is scarce or expensive to obtain. The physical states residual learning framework could potentially benefit other high-dimensional regression problems in scientific machine learning, especially when dealing with physically governed systems that exhibit stability properties.
2 Background and Related Work
2.1 Direct Neural Operator Learning
Neural operator learning aims to learn mappings between infinite-dimensional function spaces for solving parametric PDEs (kovachki2021neural, ). The operator mapping $\mathcal{G}:\mathcal{A}\longrightarrow\mathcal{U}$ maps between Banach spaces $\mathcal{A}$ and $\mathcal{U}$. For steady-state problems like Darcy Flow, the input function $a(x)\in\mathcal{A}$ typically represents coefficients while $u(x)\in\mathcal{U}$ is the solution. For time-series problems like Navier-Stokes, $a(x)$ commonly represents prior states, and $u(x)$ represents future states (li2020fourier, ).
Neural operators are constructed using various network architectures, with two prominent families being kernel-based operators and DeepONet-style operators. KernelIntegral-form architectures, such as Graph Neural Operators (GNO) (kovachki2021neural, ), are often based on iterative kernel integration. Deep Operator Networks (DeepONet) (lu2019deeponet, ) instead use separate branch and trunk networks to approximate the operator. These two general forms can be represented as:
KernelIntegral-form: $\quad\mathcal{G}_{\theta}=Q\circ\sigma(W_{l}+\mathcal{K}_{l})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1})\circ P$ (1)
DeepOnet-form: $\quad\mathcal{G}_{\theta}(a)(y)=\sum_{k=1}^{p}B_{k}(a)T_{k}(y)$
In the KernelIntegral-form, $P$ and $Q$ are projection layers, $\mathcal{K}_{i}$ are learnable kernel integral operators, $W_{i}$ are local linear operators, and $\sigma$ is a nonlinear activation. In the DeepOnet-form, $B_{k}(a)$ represents the branch network outputs encoding the input function $a$, and $T_{k}(y)$ represents the trunk network outputs encoding the output coordinates $y$. It is noted that the framework proposed in this work is applicable to both forms of architectures.
@@ 2.1 Direct Neural Operator Learning
The components of these architectures are differently instantiated in previous works. Fourier Neural Operator (FNO) (li2020fourier, ) utilizes the Fourier Transform based integral operation to efficiently learn the physical dynamics using limited training data. Factorized Fourier Neural Operator (FFNO) (tran2021factorized, ) factorize the Fourier Transform along each dimension, reducing the number of network weights. Clifford Fourier Neural Operator (CFNO) (brandstetter2022clifford, ) employs the Clifford Algebra in the network architecture, incorporating geometry prior between multi-physical fields. The various neural operators could be simply integrated with the proposed Physical States Residual Learning.
@@
Other works investigate architecture improvements for different needs, including chaotic systems modeling (li2021learning, ), physical-informed instance-wise finetuning (li2021physics, ), accelerating computation (poli2022transform, ; white2023speeding, ; rahman2022u, ; xiao2025quantum, ; DBLP:journals/jcphy/GaoW23, ), spherical fields processing (bonev2023spherical, ), irregular fields learning (li2022fourier, ; hao2023gnot, ), non-periodic boundary fields modeling (liu2023spfno, ), large-scale pretraining (grady2023model, ; subramanian2023towards, ), etc. In addition, some works explore other network backbones, e.g. improved frequency-spatial domain transformation (yueholistic, ; gupta2021multiwavelet, ; zhao2022incremental, ; cao2023lno, ; tripura2023wavelet, ; wang2024latent, ; gaodynamic, ; gao2025discretization, ), convolutions (raonic2023convolutional, ), graph neural network (brandstetter2022message, ; li2020neural, ; li2020multipole, ), attention mechanism (cao2021choose, ; hao2023gnot, ; prasthofer2022variable, ; wu2024transolver, ), and diffusion models (lim2023score, ; zhou2024unisolver, ; shan2025red, ; shan2024pird, ).
Despite theoretical approximation capabilities, practical limitations persist in generalization across resolutions (li2021physics, ), handling data scarcity (vinuesa2022emerging, ), and managing biased training distributions (bonnet2022airfrans, ). Our work addresses these challenges through residual learning between physical trajectories.
2.2 Residual Learning in Previous Works
@@
Some works (yang2023resmem, ; xu2023sequential, ; zhang2018residual, ; freund1995desicion, ) learn residual between different data samples. Similar to ResMem (yang2023resmem, ), we also utilize the training set as auxiliary memory during inference, enjoying its generalization enhancement property. The difference lies that (1) we memorize the auxiliary labels and learn the residuals, while ResMem memorizes the residuals and learns the labels, (2) our framework is end-to-end trained while ResMem is trained in separate stages. Similar external memory augmented strategies are popular in the Language Modeling community (wang2022training, ; guu2020retrieval, ).
@@
Other works learn residuals between low resolution and high resolution representation for image super-resolution (jessica2023finite, ) and deep learning based CFD simulation (jessica2023finite, ). Similar residual learning is investigated in time series prediction task  (angelopoulos2023conformal, ; xu2023sequential, ). The primary distinction between these works to ours is that we do not utilize the corresponding low-resolution solutions or previous time-step solutions, but instead, learn the residuals between different trajectories.
Additionally, (cheng2024reference, ) studies learning residuals between PDE solutions with similar geometric domains. In contrast, our work focuses on a fundamentally different problem of learning residuals between similar physical states, enabling broader applicability. We establish this based on physical system stability and develop implicit data augmentation through similarity-based sampling during training, effectively addressing distribution bias and overfitting issues in data-limited scenarios.
@@ 3 Methodology
3.1 Preliminary: Stability of Physical Systems
Stability of PDEs. For a well-posed PDE system with solution operator $\mathcal{G}$, stability means:
$\|\mathcal{G}(a_{1})-\mathcal{G}(a_{2})\|_{\mathcal{U}}\leq C\|a_{1}-a_{2}\|_{\mathcal{A}},$ (2)
where $a_{1},a_{2}$ are input functions and $C$ is a positive constant. This property, also known as Lipschitz continuity of the solution operator, is one of Hadamard’s criteria for well-posedness of PDEs. This stability property is well-established in various PDEs: elliptic PDEs satisfy stability under appropriate boundary conditions; parabolic PDEs exhibit stability through their dissipative mechanisms; and even certain chaotic systems maintain stability for suitable time intervals.
It’s noted that this stability condition is not an additional assumption imposed by our method, but rather a fundamental prerequisite for any data-driven operator learning approach to succeed (lu2019deeponet, ; li2020fourier, ). If small changes in the input function could lead to arbitrarily large, discontinuous changes in the solution, learning a generalizable mapping would be extremely challenging for neural networks.
Foundation of Residual Learning. Stability of PDEs establishes a relationship where the magnitude range of difference between output functions systematically changes with respect to the difference between input functions. This relationship allows us to explicitly control the diversity of output function residuals by selecting input function pairs with varying similarity ranges. This establishes the foundation of residual neural operator learning and enables the effectiveness of implicit data augmentation, introduced in Section˜3.5.
3.2 Residual Operator Mapping: A Unified Formulation
@@
Figure 1: The overall architecture of Physical States Residual Learning. Given an input function $a_{i}$, we first sample a similar auxiliary sample $(a_{k_{i}},u_{k_{i}})$ from the training sample set $\mathcal{T}$. Subsequently, $a_{i}$ and $(a_{k_{i}},u_{k_{i}})$ are concatenated and fed into the the neural operator $\mathcal{G}_{\theta}^{\Delta}$, producing the predicted states residual. Finally, the predicted solution $\hat{u}_{i}$ is obtained by adding the predicted states residual with the auxiliary solution $u_{k_{i}}$.
@@
Residual Operator Mapping. Based on the stability of physical systems, we formulate the residual operator mapping that captures the differences between pairs of physical trajectories:
@@
$\mathcal{A}^{2}$ $:\{(a_{i},a_{j})\mid a_{i}\in\mathcal{A},a_{j}\in\mathcal{A}\},$
$\Delta\,\mathcal{U}$ $:\{u_{i}-u_{j}\mid u_{i}\in\mathcal{U},u_{j}\in\mathcal{U}\},$
@@
Solving PDEs by Residual Operator $\mathcal{G}^{\Delta}$. In PDE solving, given an input function $a_{i}$, we first retrieve an auxiliary sample $(a_{k_{i}},u_{k_{i}})$ from the set of physical samples with known solution e.g. training set. Then the residual neural operator $\mathcal{G}^{\Delta}$ predicts the solution residual between $u_{i}$ and $u_{k_{i}}$. The final solution $u_{i}$ could be obtained by adding this predicted residual to $u_{k_{i}}$:
$u_{i}=\mathcal{G}^{\Delta}(a_{i},a_{k_{i}})+u_{k_{i}}.$ (4)
We introduce how to obtain auxiliary samples in Section˜3.3, and the specific implementation of residual neural operators in Section˜3.4.
3.3 Auxiliary Sample Retriever $\mathcal{F}^{\text{ret}}$
Given an input function $a_{i}$ and the training set $\mathcal{T}=\{a_{i},u_{i}\}_{i=1}^{N}$, we could retrieve an auxiliary sample $(a_{k_{i}},u_{k_{i}})$ from $\mathcal{T}$, formulated as follows:
$k_{i}=\mathcal{F}^{\text{ret}}(a_{i},\mathcal{T}),\,k_{i}\in\{1,2,...,N\},\,k_{i}\neq i,$ (5)
where $\mathcal{F}^{\text{ret}}$ is the sampling function. Below we present our implementation of $\mathcal{F}^{\text{ret}}$ in detail.
Training stage. During training, given an input function $a_{i}$, we randomly retrieve a sample $a_{j}$ from training set with top-k similarity scores to $a_{i}$:
$\mathcal{F}^{\text{ret}}(a_{i},\mathcal{T})$ $=\texttt{RandSample}(\texttt{argtopK}_{j,j\neq i}\,\texttt{Sim}(a_{i},a_{j})),$ (6)
$\texttt{Sim}(a_{i},a_{j})={a_{i}\cdot a_{j}\over\|a_{i}\|\|a_{j}\|},$